please tell me how to identify number of sigma and pi bonds in a hydrocarbon quickly and easily.

Grade:12

1 Answers

Bhavya
8 years ago
Here there is a trick to calculate the no of sigma and pi bonds.

Calculation of π bonds :

In the first case, we have to count the number of carbon atoms (X) and the number of hydrogen atoms (Y) in a given unsaturated hydrocarbon containing triple bonds. The formula to calculate the number of π bonds for an aliphatic open chain alkyne, where there is one or more than one triple bonds is

Where, X = number of carbon atoms; Y = number of hydrogen atoms and P = number of π bonds.

E.g.: In C16H30, X = 16, Y = 30, therefore P = [{(2X-Y)/2} + 1] = [{(2 x 16 – 30)/2} +1] = 1 + 1 = 2 number of π bonds.

Calculation of σ-bonds (S)

In the first case, we have to count the number of carbon atoms (X) and the number of hydrogen atoms (Y) in a given unsaturated hydrocarbon containing triple bonds. The formula to calculate the number of σ bonds for an aliphatic open chain alkyne, where there is one or more than one triple bonds is

Where, X = number of carbon atoms; Y = number of hydrogen atoms and S = number of σ bonds.

E.g.: In C16H30, X = 16, Y = 30, therefore, S= [X+Y-1] = [16+30-1] = 45 numbers of σ bonds.